OpenStudy (anonymous):

How would you solve this equation 4h+7h-16=6

OpenStudy (anonymous):

4h+7h-16=6 6+16=4h+7h 22=11h 2=h

OpenStudy (anonymous):

WHy do you move the 16 to the other side?

OpenStudy (anonymous):

to solve you gott to move the numbers you got to work it out and get the answer

OpenStudy (anonymous):

2 is right

OpenStudy (anonymous):

But why would you change the -16 to positive 16 and then put it adding the six?

OpenStudy (anonymous):

Because in fact, what he did was adding 16 to both sides of the equation.